OpenStudy (anonymous):

The equation x2 + y2 − 2x + 2y − 1 = 0 is the general form of the equation of a circle. What is the standard form of the equation?

OpenStudy (bradely):

x^2 + y^2 − 2x + 2y − 1 = 0 x^2 − 2x+1-1+ y^2 + 2y+1-1 − 1=0 (x-1)^2+(y+1)^2-3=0 (x-1)^2+(y+1)^2=3 (x-1)^2+(y-(-1))^2=(sqrt3)^2 Source: http://www.mathskey.com/question2answer/